Integrals Giving Inverse Trigonometric Functions


∫ dt/√(9t²-16)


1
Expert's answer
2021-04-15T07:20:41-0400

We have given the integral,

"I = \int \dfrac{dt}{\sqrt{9t^2-16}}"


"I=\int \dfrac{dt}{\sqrt{-16(1-\dfrac{9t^2}{16})}}"


"I = \int \dfrac{dt}{4i\sqrt{1-(\dfrac{3t}{4})^2}}"


Substituting "x = \dfrac{3t}{4}"

then, "dx = \dfrac{3}{4}dt"


Hence, "I = \dfrac{1}{3i}\int \dfrac{dx}{\sqrt{1-x^2}}"

Now substituting,


"x= sinu\\
dx = cosudu"


then, "I = \dfrac{1}{3i}\int\dfrac{cosu}{cosu}du"


"I = \dfrac{1}{3i}\int du"


"I = \dfrac{1}{3i}[u]+C"


Now putting the value of u in above equation


"I= \dfrac{1}{3i}sin^{-1}x + C"


Now putting the value of x in the above equation,


"I = \dfrac{1}{3i}sin^{-1}(\dfrac{3t}{4})+C"